From GlueXWiki
Jump to: navigation, search

February 7, 2013 FDC meeting


  1. Production Construction Tracking (Dave)
    • Package #2 testing (Lubomir)
  2. Basis of Estimate for the FDC activities cabling manhours (Lubomir, Bill, Dave)
  3. Corrosion tests at MePhi latest results (Lubomir)
  4. Engineering [1], [2], [3], [4], [5], [6] (Bill)
  5. Electronics (Nick)
  6. Tests at EEL126 FDC E-log (Beni)
  7. Other


Participants: Bill, Dave, Chris, Nick, Simon, Beni, and Lubomir.


- Dave for the spare parts: some issues with tensioning mylar on type-2 cathode: bubbles appeared and mylar has to be replaced. Tina will come again this week(end?) to finish the cutting of all the foils needed for the spare cathodes. Casey and Dave will string the last wire frame. On Monday Chris will put the components on the second last wire frame.

- Lubomir for package #2: despite replacing sense wire 37 on W010 (second cell), it showed a noise this time very close to the HV end (Down strip #15, Upper strip #3). On Monday we opened the package for a fourth time. During this operation, an O-ring with Apizon fell at the periphery of the top cell wire plane (W012) close to the signal side in the region of wire #76-89. Casey cleaned it with alcohol and later we checked this place with a source: no difference in the signals found. We put W010 on the stringing table and looked with the web-camera at the "noisy" place. There were some balls on the one of the field wires next to sense 37 and we cleaned it (applying some pressure with q-tips on both sides of the wire). We installed w012 in the second cell and put w010 on the top, just in case we have to open it again. This time the noise disappeared; all the channels that were re-installed were tested again and all are working fine. Next week we will continue with the ground installation.

- Additional parts to build one spare package: we realized we can re-use two spacer rings from the full-scale prototype and don't have to order them ($1.6k). Bill will look at the drawings to confirm this. They don't have guage holes but can be drilled. Bill got a quotation for the two 35mm spacer rings (to test the helium drift chamber concept) at about the same $1.6k. We will submit one PR for all the parts.

- Started testing some of the very old wire frames (W003 and W006) in a separate set-up with two Lexan sheets and a spacer ring. We need just one of them to be holding HV to have 6 spares.

Basis of Estimate for the FDC

- New version of the BOE is linked above with the breakdown of the man-weeks. During the discussions Bill added for some activities the engineer-weeks.

Corrosion test at MePhi

- Already 10 days Viton and Viton+Apiezon samples are OK, while the EPDMs failed on the first day. Vlad wants to keep the test running till the two Viton+Apiezon samples die.


- Bill is working on the procedure to cut the cooling tubes and install the manifolds there. He is considering using non-conductive tube connection at some places to avoid closed current loops. Some modifications of the cooling loops may be needed for cells 5 and 6.

- Bill showed more pictures (linked above) for the test of the cooling system. The best place to do the test is at 126; need some kind of water chiller.

Tests at EEL126

- On Tuesday we moved the fourth package from EEL to Blue Crab, took the strong-back and hubs, installed them the first package that we transported to EEL for testing. The gas is OK and after the meeting HV was set on all the package successfully.


- Dave is setting a place at the ARC building to work on the gas system control using the Brooks flow controllers.